# Learning Policies for Dynamic Coalition Formation in Multi-Robot Task Allocation

### v2

@@ Abstract
We propose a decentralized, learning-based framework for dynamic coalition formation in Multi-Robot Task Allocation (MRTA). Our approach extends Multi-Agent Proximal Policy Optimization (MAPPO) by integrating spatial action maps, robot motion planning, intention sharing, and task allocation revision to enable effective and adaptive coalition formation. Extensive simulation studies confirm the effectiveness of our model, enabling each robot to rely solely on local information to learn timely revisions of task selections and form coalitions with other robots to complete collaborative tasks. Additionally, our model significantly outperforms existing methods, including a market-based baseline. Furthermore, we evaluate the scalability and generalizability of the proposed framework, highlighting its ability to handle large robot populations and adapt to scenarios featuring diverse task sets.
@@ I Introduction
Coalition formation, widely observed in nature, such as in ant colonies and beehives, inspires researchers and engineers to design simple robots capable of strategic cooperation. In multi-robot systems, coalitions emerge when the collective actions of robots exceed what they could accomplish individually, as illustrated in Fig. 1. In this work, we focus on developing policies for a team of robots capable of performing tasks that require coalition in dynamic environments. In such scenarios, robots must learn to coordinate by leveraging local information gathered from their own sensors and shared data from neighboring robots to effectively execute multi-robot tasks.
@@
Figure 1: The firetruck dispatch as an illustrative example of a dynamic coalition formation problem: a team of firetrucks must form coalitions to tackle tasks, such as extinguishing fires, that cannot be accomplished individually.
We focus on developing a decentralized solution where each robot observes only its immediate surroundings. With such partial observability in MRTA, robots only identify tasks within their sensor range, causing the set of perceived tasks to change as they move. Additionally, we consider dynamic environments where new tasks continuously emerge over time. Well-established approaches to partially observable task allocation problems rely on inter-robot communication, where robots share their observations, internal state, and decision with their neighbors. Market-based approaches [5, 6] are among the most popular frameworks due to their robustness and scalability. However, these methods often rely on high-frequency communication for information exchange between robots to reach agreement before task assignments are made [6, 7, 8]. Although this strategy is efficient in some restricted cases, it becomes impractical in dynamic environments.
@@
    To the best of our knowledge, this is the first work to tackle the problem of decentralized dynamic coalition formation under partial observability. To facilitate coalition, we extend a Multi-Agent Reinforcement Learning (MARL) algorithm by integrating receding horizon task planning and communication of an intention map, which captures the information on the task execution plan.
@@
    Each robot’s policy is modeled as an end-to-end convolutional neural network based on the U-Net architecture [9]. The network processes information on nearby tasks and other robots within the robot’s sensing range and integrates the intentions of neighboring robots. It is trained using the MAPPO algorithm with the parameter-sharing technique.
@@
    Through extensive simulations, we demonstrate that our approach outperforms existing methods including a market-based approach, Project Manager-oriented Coalition Formation Algorithm (PCFA). Additionally, we demonstrate the scalability of models up to 1000 robots and their generalizability across a wide range of task settings.
@@ II Related Work
Solutions to ST-MR-TA MRTA include optimization-based [10, 11, 12, 13], trait-based [14, 15], and market-based [6, 8, 16, 7] methods. Among decentralized solutions for MRTA, market-based approaches are particularly popular. Specifically, PCFA [8] is a market-based solution designed for coalition formation in homogeneous teams with dynamic tasks. Due to the similarity in problem formulation, we selected PCFA as the baseline for our simulations reported in §V.
@@
Reinforcement Learning (RL) focuses on learning a policy (i.e., a mapping from observations to actions) that maximizes a robot’s total reward in a Markov decision process. It has been widely applied to various robotics problems, particularly in object manipulation [17, 18, 19, 20, 21]. While many approaches aim to directly control robot actuators, studies such as [17, 18, 22] propose the use of an abstract action map. In this framework, the policy selects a task at each time step, while a motion planner handles the low-level actuator control to navigate the robot to the task location. This abstraction alleviates the learned policy from handling low-level control, allowing it to concentrate on long-term planning.
@@
This approach is extended to the multi-robot domain in [23], where robots not only select task locations but also communicate their chosen task locations to nearby robots. The authors show that this additional communication improves cooperation and enhances overall performance. However, their method does not address coalition formation. As demonstrated in §V, their approach, with appropriate adaptations, performs suboptimally compared to ours in scenarios that require effective coalition formation. MARL naturally extends the conventional RL framework to cooperative multi-robot scenarios, where instead of learning policies to maximize individual rewards, robots learn to maximize a global reward through coordination [24]. Several studies have adapted RL algorithms for the multi-robot domain [25, 26, 27, 28, 29, 30, 31, 32, 33]. A widely adopted approach is the Centralized Training and Decentralized Execution (CTDE) paradigm, where robots are trained using shared information [27, 28, 29, 30, 31, 32, 33]. CTDE offers faster training and addresses non-stationarity, a significant challenge in MARL.
@@
Several studies have applied MARL to MRTA [29, 34, 35, 23]. For example, [29, 34] specifically address MRTA problems involving multi-robot tasks. The MRTA problem was introduced to the MARL community as a benchmark in [29], where the authors released a Python package called Level-Based Foraging (LBF) [29]. In LBF, each object is assigned a level, and it can only be picked up by a group of robots if the sum of the robots’ levels meets or exceeds the object’s level. Although LBF has become a popular MARL benchmark [36, 37, 38], it lacks support for dynamic object spawning and abstract action spaces (e.g., selecting task positions on the grid). In [36], several MARL algorithms were evaluated across popular benchmarks, including LBF (in fully observable scenarios only), and the EPyMARL package was released for community use. We expand their implementation by incorporating robot communication and additional features, which are detailed further in §IV.
@@ III Problem Description
We consider a rectangular environment represented as a 2D grid with dimensions $\left(W\times W\right)$, bounded by walls on all sides, where robots and tasks are distributed, as illustrated in Fig. 2. Each cell $p$ in the grid belongs to the cell space =$\mathbb{P}=\{1,\dots,W\}\times\{1,\dots,W\}$ and can be either empty or occupied by a robot, a task, or a wall. The grid contains a homogeneous team of $N$ robots, where the position of the $i$-th robot at time step $t$ is denoted as $p_{i,t}\in\mathbb{P}$. At each $t$, a robot can either remain in its current position or move to an adjacent cell, including diagonal moves. Robots observe the cells within their view range, defined by $R_{\text{view}}$ cells, and communicate with others within a communication range of $R_{\text{comm}}$ cells.
@@
We consider two task spawn settings: (i) Bernoulli ($p$) spawn: At each time step $t$, at each unoccupied cell, a new task can appear with probability $p$. The level of each newly spawned task is sampled from a discrete uniform distribution $\mathcal{U}(\{1,\dots,L_{max}\})$, ensuring equal likelihood for all task levels. (ii) Instant respawn: When a task is completed, a new task of the same level immediately respawns in one of the unoccupied cells. This setting maintains a constant task density throughout the episode.
To design a decentralized policy under partial observability, we model the problem as a Decentralized Partially-Observable Markov Decision Process (Dec-POMDP) [39, 40]. A Dec-POMDP is defined by the tuple $(N,\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{O},O,R)$, where $N$ is the number of robots, $\mathcal{S}$ represents the state space of the environment, and $\mathcal{A}$ and $\mathcal{O}$ denote the action and observation spaces of individual robots, respectively. The state transition function $\mathcal{T}:\mathcal{S}\times\mathcal{S}\times\mathcal{A}^{N}\to\mathbb{R}_{+}$ defines the probability $\mathcal{T}(s_{t+1}|s_{t},\mathbf{a}_{t})$ of transitioning to the next state $s_{t+1}\in\mathcal{S}$, given the current state $s_{t}\in\mathcal{S}$ and the joint action $\mathbf{a}_{t}=(a_{1,t},\cdots,a_{N,t})\in\mathcal{A}^{N}$ of all robots. The observation function $O:\mathcal{S}\times\{1,\dots,N\}\to\mathcal{O}$ provides each robot $i$ with an observation $o_{i,t}=O(s_{t},i)$, mapping the state space $\mathcal{S}$ and robot index $i$ to the observation space $\mathcal{O}$. Lastly, the shared reward function is represented by $R:\mathcal{S}\times\mathcal{A}^{N}\to\mathbb{R}$.
Each robot $i$ independently selects an action $a_{i,t}$ based on its policy $\pi:\mathcal{O}\times\mathcal{A}\to\mathbb{R}_{+}$. State-action trajectories in a Dec-POMDP are typically represented as an episode $\xi:=(s_{0},\mathbf{a}_{0},s_{1},\mathbf{a}_{1},s_{2},\mathbf{a}_{2},\dots,s_{T})$. Given an initial state distribution $P_{0}:\mathcal{S}\to\mathbb{R}_{+}$ and the episode likelihood $P_{\pi}(\xi)=P_{0}(s_{0})\prod_{t=0}^{T-1}\prod_{i=0}^{N-1}\pi(a_{i,t}|o_{i,t})\,\mathcal{T}(s_{t+1}|s_{t},\mathbf{a}_{t})$, the performance of a policy $\pi$ can be evaluated using the expected episodic reward $\mathbb{E}_{\xi\sim P_{\pi}(\cdot)}\big{[}{\sum_{t=0}^{T}\gamma^{t}R(s_{t},\mathbf{a}_{t})}\big{]}$, where $\gamma$ is the discount factor and $T$ is the terminal time.
@@ IV-A2 State and Observation Spaces
The observation $O(s_{t},i)$ for robot $i$ is derived by projecting and cropping $s_{t}$ onto areas centered around the robot’s position, determined by its view range $R_{\text{view}}$ and communication range $R_{\text{comm}}$. Specifically, the robot intention channel $\mathcal{C}^{\text{intent}}_{t}$ includes $\mathcal{C}^{\text{intent}}_{j,t}$ from robot $j$ if $j$ is within $R_{\text{comm}}$ cells of robot $i$. Meanwhile, the other channels, $\mathcal{C}^{\text{robot}}_{t},\mathcal{C}^{\text{task}}_{t},\mathcal{C}^{\text{obs}}_{t}$, are obtained by cropping $s_{t}$ according to $R_{\text{view}}$.
@@ IV-B Optimal Task Allocation Policy
IV-B1 Multi-Agent Proximal Policy Optimization (MAPPO)
To compute the optimal policy, we adopt MAPPO [27], a CTDE algorithm designed for MARL. MAPPO enables robots to share experiences during the learning phase, which is particularly beneficial for training policies in teams of homogeneous robots. MAPPO is an Actor-Critic algorithm comprising two distinct networks: the actor $\pi_{\phi}$ and the critic $V_{\theta}$. The actor outputs a distribution over actions based on each robot’s observations, formally defined by the policy $\pi_{\phi}:\mathcal{O}\times\mathcal{A}\to\mathbb{R}_{+}$, where $\phi$ represents the parameters of the policy network. These parameters $\phi$ are shared across all robots, a technique known as parameter sharing [41], which improves training efficiency, especially in cooperative scenarios involving homogeneous robots. The critic computes an approximate value function $V_{\theta}:\mathcal{S}\to\mathbb{R}$, parameterized by $\theta$, to estimate the expected discounted sum of future rewards $\mathbb{E}\big{[}{\sum_{\tau=t}^{T}\gamma^{\tau-t}R(s_{\tau},\mathbf{a}_{\tau})}\big{]}$.
@@ IV-B2 Learning Task Allocation and Revision
As illustrated in Fig. 2, we design a task allocation policy that assigns a task location $a_{i,t}$ to each robot based on its local information: $a_{i,t}\sim\pi_{\phi}(\cdot|o_{i,t})$, where $o_{i,t}=O(s_{t},i)$. In a partially observable environment with dynamically spawning tasks, a robot may encounter new tasks and other robots as it moves, which were previously unaccounted for. This dynamic nature makes it challenging for multiple robots to form coalitions for higher-level tasks. Consequently, the policy must allow robots to revise their task execution plans when necessary. However, frequent task revisions can disrupt coalition formation and hinder the completion of higher-level tasks.
@@ V-A Simulation Setup and Performance Metric
In the homogeneous scenario, tasks are evenly distributed across the grid, maintaining a uniform density. Conversely, the non-homogeneous scenario features an uneven distribution of tasks, leading to variations in task density. To define non-homogeneous configurations, we adopt the random corner patch distribution: the grid is divided into four equal-sized corner patches, and at the start of each episode, one patch is randomly chosen to contain all tasks. Within the chosen patch, the task density is uniform, while the rest of the grid remains empty. The non-homogeneous scenarios are used to evaluate whether the learned policy enables robots to form effective coalitions for task execution while efficiently locating tasks.

### v3

@@ Abstract
We propose a decentralized, learning-based framework for dynamic coalition formation in Multi-Robot Task Allocation (MRTA). Our approach extends MAPPO by integrating spatial action maps, robot motion planning, intention sharing, and task allocation revision to enable effective and adaptive coalition formation. Extensive simulation studies confirm the effectiveness of our model, enabling each robot to rely solely on local information to learn timely revisions of task selections and form coalitions with other robots to complete collaborative tasks. The results also highlight the proposed framework’s ability to handle large robot populations and adapt to scenarios with diverse task sets.
@@ I Introduction
Coalition formation, widely observed in nature, such as in ant colonies and beehives, inspires researchers and engineers to design simple robots capable of strategic cooperation. In multi-robot systems, coalitions emerge when the collective actions of robots exceed what they could accomplish individually, as illustrated in Fig. 1. In this work, we focus on developing policies for a team of robots capable of performing tasks that require coalition in dynamic environments. In such scenarios, robots must learn to coordinate by leveraging local information gathered from their own sensors and shared data from neighboring robots to effectively execute tasks.
@@
Figure 1: The firetruck dispatch as an illustrative example of a dynamic coalition formation problem: a team of firetrucks must form coalitions to extinguish fires that cannot be accomplished individually. In the event of a catastrophe that disrupts communication infrastructure, firetrucks may need to independently determine which fires to prioritize.
We focus on developing a decentralized solution where each robot observes only its immediate surroundings. With such partial observability in MRTA, robots only identify tasks within their sensor range, causing the set of perceived tasks to change as they move. Additionally, we consider dynamic environments where new tasks continuously emerge over time. Well-established approaches to partially observable task allocation problems rely on inter-robot communication, where robots share their observations, internal state, and decision with their neighbors. Market-based approaches [5, 6] are among the most popular frameworks due to their robustness and scalability. However, these methods often rely on high-frequency communication between robots to reach agreement before task assignments are made [6, 7, 8]. Although this strategy is efficient in some restricted cases, it becomes impractical in dynamic environments.
@@
    To the best of our knowledge and based on a recent survey article [9], the problem of decentralized dynamic coalition formation under partial observability has not been previously addressed. To facilitate coalition, we extend a Multi-Agent Reinforcement Learning (MARL) algorithm by integrating spatial action maps, robot motion planning, intention sharing, and task allocation revision. Compared to existing MARL approaches, such as [10, 11, 12], the integration of these key components fosters coalition formation through decentralized decision-making in dynamic environments (see Table II(c) for a comprehensive comparison).
@@
    Each robot’s policy is modeled as an end-to-end convolutional neural network based on the U-Net architecture [13]. The network processes information on nearby tasks and other robots within the robot’s sensing range and integrates the intentions of neighboring robots. It is trained using the Multi-Agent Proximal Policy Optimization (MAPPO) algorithm with the parameter-sharing technique.
@@
    Through extensive simulations, we demonstrate that our approach outperforms existing methods including a market-based approach, Project manager-oriented Coalition Formation Algorithm (PCFA). We also demonstrate its scalability with up to 1000 robots and its generalizability across a wide range of task settings.
@@ II Related Work
MRTA is widely researched topic at the intersection of robotics and multi-agent systems [14, 15, 16, 17, 18, 19, 20, 21, 22, 6, 8, 23, 7]. Among the various problem settings, ST-MR-TA MRTA formulations are most relevant to our work. Solutions to ST-MR-TA MRTA include optimization-based [16, 17, 18, 19], trait-based [21, 22], and market-based [6, 8, 23, 7] methods. Among decentralized solutions for MRTA, market-based approaches are particularly popular. Specifically, PCFA [8] is a market-based solution designed for coalition formation in homogeneous teams with dynamic tasks. Due to the similarity in problem formulation, we selected PCFA as the baseline for our simulations reported in §V.
@@
Reinforcement Learning (RL) focuses on learning a policy (i.e., a mapping from observation to action) that maximizes a robot’s total reward in a Markov decision process. It has been widely applied to various robotics problems, particularly in object manipulation [11, 24, 25, 26]. While many approaches aim to directly control robot actuators, studies such as [11, 24, 27] propose the use of an abstract action map. In this framework, the policy selects a task at each time step, while a motion planner handles the low-level actuator control to navigate the robot to the task location. This abstraction alleviates the learned policy from handling low-level control, allowing it to concentrate on long-term planning.
@@
This approach is extended to the multi-robot domain in [12], where robots not only select task locations but also communicate their chosen task locations to nearby robots. The authors show that this additional communication improves cooperation and enhances overall performance. However, their method does not address coalition formation. As demonstrated in §V, their approach, with appropriate adaptations, performs suboptimally compared to ours in scenarios that require effective coalition formation. MARL naturally extends the conventional RL framework to cooperative multi-robot scenarios, where instead of learning policies to maximize individual rewards, robots learn to maximize a global reward through coordination [28]. Several studies have adapted RL algorithms for the multi-robot domain [29, 30, 31, 32]. A widely adopted approach is the Centralized Training and Decentralized Execution (CTDE) paradigm, where robots are trained using shared information [30, 31, 32]. CTDE offers faster training and addresses non-stationarity.
@@
Several studies have applied MARL to MRTA [31, 33, 10, 12]. For example, [31, 33] specifically address MRTA problems involving multi-robot tasks. The MRTA problem was introduced to the MARL community as a benchmark in [31], where the authors released a Python package called Level-Based Foraging (LBF). In LBF, each object is assigned a level, and it can only be picked up by a group of robots if the sum of the robots’ levels meets or exceeds the object’s level. Although LBF has become a popular MARL benchmark [34, 35, 36], it lacks support for dynamic object spawning and abstract action spaces (e.g., selecting task positions on the grid). In [34], several MARL algorithms were evaluated across popular benchmarks, including LBF (in fully observable scenarios only), and the EPyMARL package was released for community use. We expand their implementation by incorporating robot communication and additional features, which are detailed further in §IV.
@@ III Problem Description
Consider a rectangular 2D grid environment of size $W\times W$, bounded by walls on all sides, where robots and tasks are distributed, as illustrated in Fig. 2. Each cell $p$ in the grid belongs to the cell space =$\mathbb{P}=\{1,\dots,W\}\times\{1,\dots,W\}$ and can be either empty or occupied by a robot, a task, or a wall. The grid contains a homogeneous team of $N$ robots, where the position of the $i$-th robot at time step $t$ is denoted as $p_{i,t}\in\mathbb{P}$. At each $t$, a robot can either remain in its current position or move to an adjacent cell, including diagonal moves. Robots observe the cells within their view range, defined by $R_{\text{view}}$ cells, and communicate with others within a communication range of $R_{\text{comm}}$ cells.
@@
We consider two task spawn settings: (i) Bernoulli ($p$) spawn: At each time step $t$, a new task can appear in any unoccupied cell with probability $p$. The level of each newly spawned task is sampled from a discrete uniform distribution $\mathcal{U}(\{1,\dots,L_{max}\})$. (ii) Instant respawn: When a task is completed, a new task of the same level immediately respawns in one of the unoccupied cells. This setting maintains a constant task density throughout the episode.
To design a decentralized policy under partial observability, we model the problem as a Decentralized Partially-Observable Markov Decision Process (Dec-POMDP) [37, 38]. A Dec-POMDP is defined by the tuple $(N,\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{O},O,R)$, where $N$ is the number of robots, $\mathcal{S}$ represents the state space of the environment, $\mathcal{A}$ and $\mathcal{O}$ denote the action and observation spaces of individual robots, respectively. The state transition function is defined as $\mathcal{T}:\mathcal{S}\times\mathcal{S}\times\mathcal{A}^{N}\to\mathbb{R}_{+}$, and the reward function as $R:\mathcal{S}\times\mathcal{A}^{N}\to\mathbb{R}$ [^1] Here, we provide a general description of a Dec-POMDP; implementation details specific to our framework are presented in §IV-A.
Each robot $i$ receives an observation $o_{i,t}\in\mathcal{O}$ at time $t$ and independently selects an action $a_{i,t}\in\mathcal{A}$ based on its (stochastic) policy $\pi:\mathcal{A}\times\mathcal{O}\to\mathbb{R}_{+}$, i.e., $a_{i,t}\sim\pi(\cdot|o_{i,t})$. An observation $o_{i,t}$ is a function of the state $s_{t}\in\mathcal{S}$ and is defined as $o_{i,t}=O(s_{t},i)$, where $O:\mathcal{S}\times\{1,\dots,N\}\to\mathcal{O}$ is the observation function. Given the state $s_{t}$ and the joint action $\mathbf{a}_{t}=(a_{1,t},\dots,a_{N,t})\in\mathcal{A}^{N}$, the next state $s_{t+1}\in\mathcal{S}$ is sampled according to the state transition function $\mathcal{T}$, i.e., $s_{t+1}\sim\mathcal{T}(\cdot|s_{t},\mathbf{a}_{t})$. The transition function $\mathcal{T}$ captures both deterministic dynamics (e.g., robot movement, task execution) and stochastic dynamics (e.g., task spawning) of the environment. For each state-action pair, a shared reward $r_{t}=R(s_{t},\mathbf{a}_{t})$ is computed for all robots.
In a Dec-POMDP, state-action trajectories are represented as an episode $\xi:=(s_{0},\mathbf{a}_{0},s_{1},\mathbf{a}_{1},\dots,s_{T})$. Given an initial state distribution $P_{0}:\mathcal{S}\to\mathbb{R}_{+}$ and the episode likelihood $P_{\pi}(\xi)=P_{0}(s_{0})\prod_{t=0}^{T-1}\prod_{i=0}^{N-1}\pi(a_{i,t}|o_{i,t})\,\mathcal{T}(s_{t+1}|s_{t},\mathbf{a}_{t})$, the performance of a policy $\pi$ can be evaluated using the expected episodic reward $\mathbb{E}_{\xi\sim P_{\pi}(\cdot)}\big{[}{\sum_{t=0}^{T}R(s_{t},\mathbf{a}_{t})}\big{]}$, where $T$ is the terminal time. When a sufficiently large number of episodes is available in a dataset $\mathcal{D}$, this expectation can be approximated as $\mathbb{E}_{\xi\sim\mathcal{D}}[{\sum_{t=0}^{T}R(s_{t},\mathbf{a}_{t})}]$, thereby removing the need for explicit knowledge of $P_{0}$ and $\mathcal{T}$.
@@ IV-A2 State and Observation Spaces
The observation $O(s_{t},i)$ for robot $i$ is derived by projecting and cropping $s_{t}$ onto areas centered around the robot’s position, determined by its view range $R_{\text{view}}$ and communication range $R_{\text{comm}}$. Specifically, the robot intention channel $\mathcal{C}^{\text{intent}}_{t}$ includes $\mathcal{C}^{\text{intent}}_{j,t}$ from robot $j$ if $j$ is within $R_{\text{comm}}$ cells of robot $i$. The other channels, $\mathcal{C}^{\text{robot}}_{t},\mathcal{C}^{\text{task}}_{t},\mathcal{C}^{\text{obs}}_{t}$, are obtained by cropping $s_{t}$ according to $R_{\text{view}}$.
@@ IV-B Optimal Task Allocation Policy
IV-B1 MAPPO
To compute the optimal policy, we adopt MAPPO [30], a CTDE algorithm designed for MARL. MAPPO enables robots to share experiences during the learning phase, which is particularly beneficial for training policies in teams of homogeneous robots. MAPPO is an Actor-Critic algorithm comprising two distinct networks: the actor $\pi_{\phi}$ and the critic $V_{\theta}$. The actor outputs a distribution over actions based on each robot’s observations, formally defined by the policy $\pi_{\phi}:\mathcal{A}\times\mathcal{O}\to\mathbb{R}_{+}$, where $\phi$ represents the parameters of the policy network. These parameters $\phi$ are shared across all robots, a technique known as parameter sharing [39], which improves training efficiency, especially in cooperative scenarios involving homogeneous robots. The critic computes an approximate value function $V_{\theta}:\mathcal{S}\to\mathbb{R}$, parameterized by $\theta$, to estimate the expected discounted sum of future rewards $\mathbb{E}[{\sum_{\tau=t}^{T}\gamma^{\tau-t}R(s_{\tau},\mathbf{a}_{\tau})}]$.
@@ IV-B2 Learning Task Allocation and Revision
As shown in Fig. 2, we design a task allocation policy that assigns a task location $a_{i,t}$ to each robot based on its local information: $a_{i,t}\sim\pi_{\phi}(\cdot|o_{i,t})$, where $o_{i,t}=O(s_{t},i)$. In a partially observable environment with dynamically spawning tasks, a robot may encounter new tasks and other robots as it moves, which were previously unaccounted for. This dynamic nature makes it challenging for multiple robots to form coalitions for higher-level tasks. Consequently, the policy must allow robots to revise their task execution plans when necessary. However, frequent task revisions can disrupt coalition formation and hinder the completion of higher-level tasks.
@@ V-A Simulation Setup and Performance Metric
In the homogeneous scenario, tasks are evenly distributed across the grid, maintaining a uniform density. Conversely, the non-homogeneous scenario features an uneven distribution of tasks, leading to variations in task density. To define non-homogeneous configurations, we adopt the random corner patch distribution: the grid is divided into four equal-sized corner patches, and at the start of each episode, one patch is randomly chosen to contain all tasks. Within the chosen patch, the task density is uniform, while the rest of the grid remains empty. These scenarios are used to evaluate whether the learned policy enables robots to form effective coalitions for task execution while efficiently locating tasks.
@@
Computational Resources: The policies were trained for 4M steps over a period of up to 13 hours and up to 68 hours in the non-homogeneous and homogeneous settings, respectively. Each policy is trained using 16 CPU cores (Intel Xeon Platinum 8260) and one GPU (Nvidia V100). After training, however, each trained policy executes in just $2.1\pm 1.5$ ms on a single CPU core of the same model.